# Active Learning to Guide Labeling Efforts for Question Difficulty Estimation

### v1

@@ 5.2 Acquisition Behavior
As expected, Uniform acquisition causes minimal changes in the level distribution because samples are randomly selected from the unlabeled pool. In contrast, Variance acquisition exhibits a distinctive pattern, selecting many level 2 observations and few level 0 instances. The proportion of level 2 samples increases from 13% to 45%, while level 0 samples decrease from 25% to merely 6%. These findings partially align with previous studies Atighehchian et al. (2020) suggesting that top-$K$ strategies using epistemic uncertainty target underrepresented classes. Variance acquisition indeed prioritizes sampling from the most underrepresented class (level 2) but does this primarily at the expense of level 0 observations, rather than the majority class (level 1).

### v2

@@ 5.2 Acquisition Behavior
As expected, Uniform acquisition causes minimal changes in the level distribution because samples are randomly selected from the unlabeled pool. In contrast, Variance acquisition exhibits a distinctive pattern, selecting many level 2 observations and few level 0 instances. The proportion of level 2 samples increases from 13% to 45%, while level 0 samples decrease from 25% to merely 6%. These findings partially align with previous studies (Atighehchian et al., 2020) suggesting that top-$K$ strategies using epistemic uncertainty target underrepresented classes. Variance acquisition indeed prioritizes sampling from the most underrepresented class (level 2) but does this primarily at the expense of level 0 observations, rather than the majority class (level 1).